5.    Suppose that  

i. For what values of  are  and  mutually exclusive? For this value of , are

                          A and B independent?

           ii. For what values of , are  and  independent?

           Determine whether for this value of x, A and B are mutually exclusive.

 



1
Expert's answer
2021-07-19T05:51:46-0400
"8P(A\cup B)=5=>P(A\cup B)=\dfrac{5}{8}"

"\dfrac{2x}{P(A)}=1, P(B)=x"

i. If "A" and "B" are mutually exclusive, then "P(A\cap B)=0." Hence


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=P(A)+P(B)-0=P(A)+P(B)"

"\dfrac{5}{8}=2x+x"

"x=\dfrac{5}{24}"

If "A" and "B" are independent, then "P(A\cap B)=P(A)P(B)."


Since "P(A\cap B)=0, P(A)\not=0, P(B)\not=0," then "A" and "B" are not independent.


ii.

If "A" and "B" are independent, then "P(A\cap B)=P(A)P(B)"

Then


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=P(A)+P(B)-P(A)P(B)"

"\dfrac{5}{8}=2x+x-x(2x)"

"16x^2-24x+5=0"

"x=\dfrac{24\pm\sqrt{(24)^2-4(16)(5)}}{2(16)}=\dfrac{3\pm2}{4}"

Since "0\leq x\leq\dfrac{1}{2}," we take "x=\dfrac{3-2}{4}=\dfrac{1}{4}."


"P(A)=\dfrac{1}{2}, P(B)=\dfrac{1}{4}, P(A\cap B)=\dfrac{1}{2}(\dfrac{1}{4})=\dfrac{1}{8}\not=0"

Since "P(A\cap B)\not=0," then "A" and "B" are not mutually exclusive.